# Event Data Association via Robust Model Fitting for Event-based Object Tracking

### v2

@@ 4.2 Evaluation Protocol
The test object tracking sequences include the shapes_translation, shapes_6dof, poster_6dof, and slider_depth sequences from the popular Event Camera Dataset (ECD) mueggler2017event and the fast_drone, light_variations, what_is_background, and occlusions sequences from the challenging Extreme Event Dataset (EED) mitrokhin2018event . In the four sequences from the ECD dataset, the first three sequences have an increasing motion speed. The fourth sequence has a constant motion speed. For the four sequences, the corresponding object textures vary from simple Black&White shapes to complicated artifacts. By using these four sequences, we are mainly concerned with the performance of all competing methods for a variety of object motions, especially for fast motion, and for different object shapes and textures. For the four sequences from the EED dataset, the first three sequences record a fast moving drone under low illumination environments. The fourth sequence records a moving ball with a net as the foreground. By using these four sequences, we evaluate the performance of the competing trackers in low illumination conditions and occlusion situations.
For the evaluation metrics, we follow the frame-wise tracking protocol in chen2020end and use the Average Overlap Rate (AOR) and the Average Robustness (AR) to respectively measure the precision and the robustness of all the competing object tracking methods, as follows:
@@ 4.3 Ablation Study
Method shapes_translation shapes_6dof poster_6dof slider_depth
SiamBAN chen2022siamban 0.826 0.811 0.832 0.907
SiamRPN++ Li_2019_CVPR 0.790 0.779 0.753 0.907
ATOM Danelljan_2019_CVPR 0.815 0.803 0.835 0.897
ECO-E danelljan2017eco 0.821 0.834 0.783 0.771
E-MS barranco2018real 0.675 0.612 0.417 0.447
EVT messikommer2023data 0.873 0.845 0.816 0.938
ETD chen2019asynchronous 0.843 0.852 0.824 0.903
RMRNet chen2020end 0.836 0.866 0.859 0.915
EDA(Ours) 0.896 0.866 0.872 0.969
Method shapes_translation shapes_6dof poster_6dof slider_depth
SiamBAN chen2022siamban 0.945 0.967 0.954 1.000
SiamRPN++ Li_2019_CVPR 0.942 0.972 0.899 1.000
ATOM Danelljan_2019_CVPR 0.945 0.974 0.961 1.000
ECO-E danelljan2017eco 0.941 0.960 0.878 0.993
E-MS barranco2018real 0.768 0.668 0.373 0.350
EVT messikommer2023data 0.971 0.988 0.979 1.000
ETD chen2019asynchronous 0.998 0.998 0.995 1.000
RMRNet chen2020end 0.951 0.980 0.962 1.000
EDA(Ours) 0.998 0.998 0.983 1.000
SiamBAN chen2022siamban 0.764 0.771 0.676 0.148
SiamRPN++ Li_2019_CVPR 0.717 0.497 0.653 0.096
ATOM Danelljan_2019_CVPR 0.763 0.652 0.725 0.387
ECO-E danelljan2017eco 0.728 0.685 0.099 0.308
E-MS barranco2018real 0.313 0.325 0.362 0.356
EVT messikommer2023data 0.836 0.813 0.672 0.585
ETD chen2019asynchronous 0.817 0.859 0.653 0.608
RMRNet chen2020end 0.892 0.802 0.202 0.716
EDA(Ours) 0.913 0.894 0.752 0.702
SiamBAN chen2022siamban 1.000 0.852 0.841 0.167
SiamRPN++ Li_2019_CVPR 0.941 0.500 0.833 0.167
ATOM Danelljan_2019_CVPR 1.000 0.921 0.917 0.500
ECO-E danelljan2017eco 0.882 0.803 0.000 0.333
E-MS barranco2018real 0.307 0.321 0.360 0.353
EVT messikommer2023data 1.000 0.932 0.825 0.641
ETD chen2019asynchronous 0.941 0.934 0.807 0.647
RMRNet chen2020end 1.000 0.947 0.083 0.833
EDA(Ours) 1.000 0.966 0.917 0.833
@@
In addition, we perform a parameter analysis on the challenging light_ variations sequence. Firstly, we study the influence of different inlier noise scale $\tau$. The proposed EDA respectively achieves AOR/AR scores of 0.889/0.963, 0.892/0.966, 0.894/0.966, 0.894/0.966, and 0.883/0.957 when $\tau$ is set to 0.0025, 0.005, 0.01, 0.02, and 0.05. From the results, we can see that the performance of EDA is not very sensitive to the value of $\tau$, and a small value of $\tau$ (around 0.01) can work well even in challenging scenes. Then, we replace $\tau$ with the IKOSE estimator wang2011simultaneously . Since IKOSE needs the inlier ratio $k$, we study the influence of different $k$. EDA respectively obtains AOR/AR scores of 0.859/0.947, 0.883/0.966, 0.898/0.966, 0.871/0.947, and 0.847/0.941 when $k$ is set to 0.0025, 0.005, 0.01, 0.02, and 0.05. The comparison between the EDA with a fixed $\tau$ and the EDA with IKOSE are shown in the left image of Fig. 5. As shown in the figure, the EDA with IKOSE only achieves a small improvement when $k$ is set to 0.01, and its AOR/AR scores drops considerably when $k$ is set to the other values compared with the EDA with a fixed $\tau$. Therefore, the proposed EDA adopts the fixed noise scale $\tau$ to avoid the trivial parameter tuning. Finally, we study the influence of different numbers of the time slices $N_{s}$ on the performance of EDA. EDA respectively achieves AOR/AR scores of 0.853/0.931, 0.897/0.966, 0.894/0.966, 0.769/0.941, and 0.691/0.875 when $N_{s}$ is set to 3, 5, 10, 15, and 20, as shown in the right image of Fig. 5. The proposed EDA with $N_{s}$=5 only achieves 0.003 improvements on AOR compared with the EDA with $N_{s}$=10 at the cost of a 10x slower speed. Therefore, it is a good trade-off between accuracy and efficiency to set $N_{s}$ to 10.
4.4 Event-Based Object Tracking
Figure 6: Quantitative tracking results obtained by SiamBAN, EVT, SiamRPN++, ATOM, E-MS, ETD, RMRNet, and the proposed EDA. Each row represents a sequence. From top to bottom, the corresponding sequences are shape_6dof, poster_6dof, slider_depth, light_variations, what_is_background, and occlusions, respectively. From left to right, the 2nd, 4th, 6th, and 8th columns show the results obtained by ETD, RMRNet, and EDA with the ground truth (abbreviated as GT). The 1st, 3rd, 5th, and 7th columns show the results obtained by the other competing methods with the GT. Best viewed in color.
In this work, we compare the proposed EDA with eight popular tracking methods, including SiamBAN chen2022siamban , SiamRPN++ Li_2019_CVPR , ATOM Danelljan_2019_CVPR , EVT messikommer2023data , E-MS barranco2018real , ETD chen2019asynchronous , RMRNet chen2020end , and an event-based variant of the classical tracker ECO danelljan2017eco (called as ECO-E). Among these trackers, SiamBAN, SiamRPN++, and ATOM are the popular conventional tracking methods. EVT, E-MS, ETD, and RMRNet are the popular event-based tracking methods. ECO-E is an event-based variant of ECO, which uses TSLTD chen2020end event frames as its inputs. Thus, ECO-E is adopted to evaluate the performance of ECO on the event data. Moreover, E-MS and EVT are extended to support the bounding box-based object tracking.
The quantitative results are given in Table 1, Table 2, Table 3, and Table 4. We also provide some qualitative results obtained by SiamBAN, EVT, SiamRPN++, ATOM, E-MS, ETD, RMRNet, and our EDA in Fig. 6. From the tables and the figure, we can see that our EDA achieves the best performance on most of the sequences except for the poster_6dof and occlusions sequences, on which our EDA has obtained the second best AR and AOR, respectively. This is because that, these two sequences (i.e., poster_6dof and occlusions) include considerable object pairs that are almost out-of-view or highly occluded. These object pairs slightly affect the robustness of EDA compared with the edge-based ETD tracker and the motion estimation-based RMRNet tracker. The superior performance achieved by EDA is largely attributed to the proposed two-stage weighting algorithm, which accurately selects true event trajectories without being affected by the sensor noise.
As for the competitors, SiamBAN, SiamRPN++, and ATOM also achieve competitive results on some sequences of the ECD dataset. However, they show inferior tracking performance on the sequences with fast motion and low illumination conditions (such as the light_variations and occlusions sequences). They usually lose the tracked objects, due to the influence of motion blur and HDR conditions. In contrast, the five event-based competitors (i.e., ECO-E, E-MS, EVT, ETD, and RMRNet) can effectively handle most fast motions and HDR scenes, by leveraging the good properties of the event data. Nevertheless, E-MS is less effective in handling the challenges of complicated object textures and cluttered backgrounds (e.g., on the poster_6dof and slider_depth sequences), and it is sensitive to sensor noise. Compared with E-MS, the EVT, ETD and RMRNet trackers achieve very competitive results on most of the sequences. However, due to the unstable performance during the processes of the edge-based object proposal generation and motion estimation, EVT, ETD and RMRNet achieve inferior precision performance compared with the proposed EDA. Moreover, compared with the TSLTD based RMRNet tracker, ECO-E has achieved much inferior performance, which indicates that directly applying conventional tracking methods to event data is less effective.
We also provide the event trajectory results obtained by the proposed EDA on some representative sequences from the ECD and EED datasets. Furthermore, in order to evaluate the performance of EDA under more tracking scenarios with challenges, we test it on some representative sequences from the Color Event Dataset (CED) scheerlinck2019ced and the recently released EVIMO dataset mitrokhin2019ev . Here, it should be noted that there is no ground truth that is publicly available for the CED and EVIMO datasets for the task of object tracking. Therefore, we only provide the event trajectory results obtained by EDA for qualitative evaluation. The qualitative results are shown in Fig. 7.
@@ 5 Conclusions
In this paper, we propose a novel unifying event data association (EDA) approach to effectively and explicitly handle the essential event data association and event information fusion problem. The proposed EDA performs a model fitting on event data, which can asynchronously associate and fuse the event data over time for motion-related computer vision tasks, such as object tracking. Specifically, EDA presents a deterministic strategy to effectively generate spatio-temporal model hypotheses from the fused retinal events. Then, thanks to the two-stage weighting, EDA can robustly weigh and select the model hypotheses without been affected by sensor noise and irrelevant information. Furthermore, different from the previous model fitting methods, EDA proposes to adaptively determine the true models, which avoids the empirical risk. Extensive experiments on several challenging datasets demonstrate the effectiveness and superiority of the proposed EDA.
In future work, we plan to explore more delicate model fitting approaches for event association, and this approach will also be extended for a variety of motion-related computer vision tasks. Besides, optical image data can provide additional visual information and may play a positive role in a hybrid information fusion with the event data. Therefore, how to simultaneously associate the event and image data to improve the performance of event-based methods is also one of the directions we will focus on in the future.
6 Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant Nos. 62036007, U21A20514, and U22A2095), the Postdoctoral Fellowship Program of China Postdoctoral Science Foundation (Grant No. GZC20233362), the Chongqing Postdoctoral Innovative Talents Support Program (Grant No. CQBX202316), and the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515011740)

### v3

@@ 4.2 Evaluation Protocol
The test object tracking sequences include the shapes_translation, shapes_6dof, poster_6dof, and slider_depth sequences from the popular Event Camera Dataset (ECD) Mueggler et al. (2017) and the fast_drone, light_variations, what_is_background, and occlusions sequences from the challenging Extreme Event Dataset (EED) Mitrokhin et al. (2018). In the four sequences from the ECD dataset, the first three sequences have an increasing motion speed. The fourth sequence has a constant motion speed. For the four sequences, the corresponding object textures vary from simple Black&White shapes to complicated artifacts. By using these four sequences, we are mainly concerned with the performance of all competing methods for a variety of object motions, especially for fast motion, and for different object shapes and textures. For the four sequences from the EED dataset, the first three sequences record a fast moving drone under low illumination environments. The fourth sequence records a moving ball with a net as the foreground. By using these four sequences, we evaluate the performance of the competing trackers in low illumination conditions and occlusion situations.
For the evaluation metrics, we follow the frame-wise tracking protocol in Chen et al. (2020) and use the Average Overlap Rate (AOR) and the Average Robustness (AR) to respectively measure the precision and the robustness of all the competing object tracking methods, as follows:
@@
Method shapes_translation shapes_6dof poster_6dof slider_depth
SiamFC Bertinetto et al. (2016) 0.812 0.835 0.830 0.909
ECO Danelljan et al. (2017) 0.823 0.847 0.846 0.947
SiamRPN++ Li et al. (2019) 0.790 0.779 0.753 0.907
ATOM Danelljan et al. (2019) 0.815 0.803 0.835 0.897
ECO-E Danelljan et al. (2017) 0.821 0.834 0.783 0.771
E-MS Barranco et al. (2018) 0.675 0.612 0.417 0.447
ETD Chen et al. (2019) 0.843 0.852 0.824 0.903
RMRNet Chen et al. (2020) 0.836 0.866 0.859 0.915
EDA(Ours) 0.896 0.866 0.872 0.969
Method shapes_translation shapes_6dof poster_6dof slider_depth
SiamFC Bertinetto et al. (2016) 0.940 0.968 0.956 1.000
ECO Danelljan et al. (2017) 0.943 0.969 0.960 1.000
SiamRPN++ Li et al. (2019) 0.942 0.972 0.899 1.000
ATOM Danelljan et al. (2019) 0.945 0.974 0.961 1.000
ECO-E Danelljan et al. (2017) 0.941 0.960 0.878 0.993
E-MS Barranco et al. (2018) 0.768 0.668 0.373 0.350
ETD Chen et al. (2019) 0.998 0.998 0.995 1.000
RMRNet Chen et al. (2020) 0.951 0.980 0.962 1.000
EDA(Ours) 0.998 0.998 0.983 1.000
SiamFC Bertinetto et al. (2016) 0.766 0.772 0.712 0.148
ECO Danelljan et al. (2017) 0.830 0.782 0.675 0.209
SiamRPN++ Li et al. (2019) 0.717 0.497 0.653 0.096
ATOM Danelljan et al. (2019) 0.763 0.652 0.725 0.387
ECO-E Danelljan et al. (2017) 0.728 0.685 0.099 0.308
E-MS Barranco et al. (2018) 0.313 0.325 0.362 0.356
ETD Chen et al. (2019) 0.817 0.859 0.653 0.608
RMRNet Chen et al. (2020) 0.892 0.802 0.202 0.716
EDA(Ours) 0.913 0.894 0.752 0.702
SiamFC Bertinetto et al. (2016) 1.000 0.947 0.833 0.000
ECO Danelljan et al. (2017) 1.000 0.934 0.750 0.333
SiamRPN++ Li et al. (2019) 0.941 0.500 0.833 0.167
ATOM Danelljan et al. (2019) 1.000 0.921 0.917 0.500
ECO-E Danelljan et al. (2017) 0.882 0.803 0.000 0.333
E-MS Barranco et al. (2018) 0.307 0.321 0.360 0.353
ETD Chen et al. (2019) 0.941 0.934 0.807 0.647
RMRNet Chen et al. (2020) 1.000 0.947 0.083 0.833
EDA(Ours) 1.000 0.966 0.917 0.833
4.3 Event-Based Object Tracking
Figure 4: Quantitative tracking results obtained by SiamFC, ECO, SiamRPN++, ATOM, E-MS, ETD, RMRNet, and the proposed EDA. Each row represents a sequence. From top to bottom, the corresponding sequences are shape_6dof, poster_6dof, slider_depth, light_variations, what_is_background, and occlusions, respectively. From left to right, the 2nd, 4th, 6th, and 8th columns show the results obtained by ETD, RMRNet, and EDA with the ground truth (abbreviated as GT). The 1st, 3rd, 5th, and 7th columns show the results obtained by the other competing methods with the GT. Best viewed in color.
In this work, we compare the proposed EDA with eight popular tracking methods, including SiamFC Bertinetto et al. (2016), ECO Danelljan et al. (2017), SiamRPN++ Li et al. (2019), ATOM Danelljan et al. (2019), E-MS Barranco et al. (2018), ETD Chen et al. (2019), RMRNet Chen et al. (2020), and an event-based variants of ECO (called as ECO-E). Among these trackers, SiamFC, ECO, SiamRPN++, and ATOM are the popular conventional tracking methods. E-MS, ETD, and RMRNet are the state-of-the-art event-based tracking methods. ECO-E is an event-based variant of ECO, which uses TSLTD frames as its inputs. Thus, ECO-E is adopted to evaluate the performance of ECO on the event data. Moreover, E-MS is extended to support the bounding box-based object tracking.
The quantitative results are given in Table 1, Table 2, Table 3, and Table 4. We also provide some qualitative results obtained by SiamFC, ECO, SiamRPN++, ATOM, E-MS, ETD, RMRNet, and our EDA in Fig. 4. From the tables and the figure, we can see that our EDA achieves the best performance on most of the sequences except for the poster_6dof and occlusions sequences, on which our EDA has obtained the second best AR and AOR, respectively. This is because that, these two sequences (i.e., poster_6dof and occlusions) include considerable object pairs that are almost out-of-view or highly occluded. These object pairs slightly affect the robustness of EDA compared with the edge-based ETD tracker and the motion estimation-based RMRNet tracker. The superior performance achieved by EDA is largely attributed to the proposed two-stage weighting algorithm, which accurately selects true event trajectories without being affected by the sensor noise.
As for the competitors, SiamFC, ECO, SiamRPN++, and ATOM also achieve competitive results on some sequences of the ECD dataset. However, they show inferior tracking performance on the sequences with fast motion and low illumination conditions (such as the light_variations and occlusions sequences). They usually lose the tracked objects, due to the influence of motion blur and HDR conditions. In contrast, the three event-based competitors (i.e., ECO-E, E-MS, and ETD) can effectively handle fast motions and HDR scenes, by leveraging the good properties of the event data. Nevertheless, E-MS is less effective in handling the challenges of complicated object textures and cluttered backgrounds (e.g., on the poster_6dof and slider_depth sequences), and it is sensitive to sensor noise. Compared with E-MS, the ETD and RMRNet trackers achieve very competitive results on most of the sequences. However, due to the unstable performance during the processes of the edge-based object proposal generation and motion estimation, ETD and RMRNet achieve inferior precision performance compared with the proposed EDA. Moreover, compared with ECO, ECO-E has achieved much inferior performance, which indicates that directly applying conventional tracking methods to event data is less effective.
We also provide the event trajectory results obtained by the proposed EDA on some representative sequences from the ECD and EED datasets. Furthermore, in order to evaluate the performance of EDA under more tracking scenarios with challenges, we test it on some representative sequences from the Color Event Dataset (CED) Scheerlinck et al. (2019) and the recently released EVIMO dataset Mitrokhin et al. (2019). Here, it should be noted that there is no ground truth that is publicly available for the CED and EVIMO datasets for the task of object tracking. Therefore, we only provide the event trajectory results obtained by EDA for qualitative evaluation. The qualitative results are shown in Fig. 5.
5 Discussion
{adjustwidth}
[…]
@@ 5.1 Ablation Study
5.2 Parameter Analysis
@@
In addition, we perform a parameter analysis on the challenging light_ variations sequence. Firstly, we study the influence of different inlier noise scale $\tau$. The proposed EDA respectively achieves AOR/AR scores of 0.889/0.963, 0.892/0.966, 0.894/0.966, 0.894/0.966, and 0.883/0.957 when $\tau$ is set to 0.0025, 0.005, 0.01, 0.02, and 0.05. From the results, we can see that the performance of EDA is not very sensitive to the value of $\tau$, and a small value of $\tau$ (around 0.01) can work well even in challenging scenes. Then, we replace $\tau$ with the IKOSE estimator Wang et al. (2011). Since IKOSE needs the inlier ratio $k$, we study the influence of different $k$. EDA respectively obtains AOR/AR scores of 0.859/0.947, 0.883/0.966, 0.898/0.966, 0.871/0.947, and 0.847/0.941 when $k$ is set to 0.0025, 0.005, 0.01, 0.02, and 0.05. The comparison between the EDA with a fixed $\tau$ and the EDA with IKOSE are shown in the left image of Fig. 7. As shown in the figure, the EDA with IKOSE only achieves a small improvement when $k$ is set to 0.01, and its AOR/AR scores drops considerably when $k$ is set to the other values compared with the EDA with a fixed $\tau$. Therefore, the proposed EDA adopts the fixed noise scale $\tau$ to avoid the trivial parameter tuning. Finally, we study the influence of different numbers of the time slices $N_{s}$ on the performance of EDA. EDA respectively achieves AOR/AR scores of 0.853/0.931, 0.897/0.966, 0.894/0.966, 0.769/0.941, and 0.691/0.875 when $N_{s}$ is set to 3, 5, 10, 15, and 20, as shown in the right image of Fig. 7. The proposed EDA with $N_{s}$=5 only achieves 0.003 improvements on AOR compared with the EDA with $N_{s}$=10 at the cost of a 10x slower speed. Therefore, it is a good trade-off between accuracy and efficiency to set $N_{s}$ to 10.
@@ 6 Conclusions
In this paper, we propose a novel unifying event data association approach (EDA) to effectively and explicitly handle the essential event data association problem. The proposed EDA can asynchronously gather the event data over time for processing. Then, we present a deterministic model hypothesis generation strategy to effectively generate model hypotheses in the spatio-temporal domain from the gathered events. After that, we provide a two-stage weighting algorithm, which robustly weighs and selects the true models from the generated model hypotheses through multi-structural model fitting. Furthermore, we propose an adaptive model selection strategy to automatically determine the number of the true models. Finally, the selected true models are exploited to perform the final data association for remote object tracking. Extensive experiments on several challenging datasets demonstrate the effectiveness and superiority of the proposed EDA.
[…]
Conceptualization, H.C. and Y.W.; methodology, H.C.; software, H.C. and Y.W.; validation, Y.W. and Y.P.; formal analysis, H.C.; investigation, H.C.; resources, H.C.; data curation, Y.W.; writing-original draft preparation, H.C.; writing—review and editing, Y.W. and Y.P.; visualization, Y.P.; supervision, Y.W. and Y.P. All authors have read and agreed to the published version of the manuscript.
[…]
This work was supported by the China Postdoctoral Science Foundation (Grant Nos. GZC20233362, 2024MD754043, and 2024MD754039), Chongqing Postdoctoral Innovative Talents Support Program (Grant No. CQBX202316), Science and Technology Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202400648), and National Natural Science Foundation of China (Grant No. 62301103).
[…]
The original contributions presented in this study are included in the article. Further inquiries can be directed to the corresponding author.
Acknowledgements.
[…]